# A causal random synthetic turbulent velocity field: I. The underlying Gaussian structure

### v1

@@ 1 Introduction
$\mathbb{E}\left[\left(\delta_{\ell}^{\text{\tiny{long}}}\mathbf{u}^{\nu}\right)^{2}\right]\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}2|\ell|^{2}\int_{k\in\mathbb{R}}\left(2\pi k|\ell|\right)^{2}E^{\text{\tiny{E,long}}}_{\nu}(k){\mathrm{d}}k\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\nu\to 0}}\frac{\varepsilon_{\nu}}{15\nu}|\ell|^{2},$ (27)
@@ 2.2 A first comparison concerning the spatial structure of DNS and the Gaussian model
We show in Fig. 1 a comparison between the velocity field $\mathbf{u}^{\text{\tiny{DNS}},\nu}(t,\mathbf{x})$ (Fig. 1(a)) extracted from a direct numerical simulations (DNSs) of the forced Navier-Stokes equations (1) and a Gaussian model (Fig. 1(b)). Concerning the DNS, it is provided by the Johns Hopkins turbulence database (Li et al., 2008). The data set concerns a simulation of the Navier-Stokes equations in a homogeneous and isotropic situation, using a pseudo-spectral method fully de-aliased with $N=1024$ nodes in each direction, over 5028 timesteps, corresponding to roughly speaking ten turnover time scales. The Taylor-based Reynolds number is estimated to be 433, corresponding to a fully developed turbulent state. Detailed properties of the simulation are provided on their website. For our purpose, we have downloaded the three velocity components over three spatial slices : $(0,y,z)$, $(x,0,z)$ and $(x,y,0)$, at all available times. The Gaussian model $\mathbf{u}^{\text{\tiny{fGv}},\nu}(\mathbf{x})$ corresponds to a fractional Gaussian vector field, in a periodic setting, presented in subsection 3.2.2. At this stage, it is enough to state that this is the unique statistically homogeneous, isotropic and incompressible Gaussian vector field consistent with the longitudinal PSD $E_{\nu}^{\text{\tiny{E,long}}}(k)$ provided in (10). Note that the three-dimensional PSD $E_{\nu}^{\text{\tiny{E}}}(k)$ appearing in the expression of the Gaussian fractional velocity field (72) is obtained from the longitudinal one (10) using the equality provided in (9). (72). Following a suitable truncation of the Fourier modes, as described in Section 4.1, we have similarly performed a simulation of this Gaussian vector field on a periodic box of size $L_{\text{\tiny{tot}}}=2\pi$ using $N=1024$ nodes in each direction, with the remaining free parameters entering in (10) chosen to match the DNS dataset. This corresponds to $D_{2}=0.021$, $L=L_{\text{\tiny{tot}}}$ and $\eta_{d}=0.085$. The results of our comparison are collected in Fig. 1.
@@
The statistical analysis of these two vector fields is shown in panels 1(c) and 1(d). The respective longitudinal PSDs $E_{\nu}^{\text{\tiny{E,long}}}(k)$ are shown in (c). Concerning the DNS, it is obtained as the variance of the respective Fourier modes obtained by performing one-dimensional discrete Fourier transforms of the longitudinal components, say, for instance, the $x$ component of velocity along the direction $x$. The statistical sample is averaged over the three longitudinal components, spatially over the slices, and over time. Results of our estimation are displayed with an orange solid line (light gray). The estimation of the Gaussian model statistics is also performed by averaging over the three components, over space and time, as has been done for the DNS. We display the result of our estimation using a solid blue line (dark gray). As we can observe, both longitudinal PSDs superimpose in a very satisfactory manner, showing that, based on the variance of Fourier modes, these two fields are statistically indiscernible. We mention that we could have, instead of the estimate on the Gaussian model, displayed the theoretical functional form provided in (10) without noticeable difference, as expected. The result of our statistical estimation of the longitudinal structure function $S_{2,\nu}^{\text{\tiny{long}}}(\ell)$ (20) is displayed in Fig. 1(d), using an orange line (light gray) for DNS and a blue line (dark gray) for the Gaussian model, averages have been taken on the same statistical sample as the one considered for Fig. 1(c). The three ranges of scales, including decorrelation length scales of order of the size of the box $L_{\text{\tiny{tot}}}$, inertial and dissipative length scales, are reproduced in a very satisfactory manner by the model, as expected from the behaviors of the PSDs of Fig. 1(c).
@@ 2.3 The statistical temporal structure of the turbulent velocity field and the sweeping effect
From the statistical spatial structure of the velocity field formerly presented, in large part given by the scaling relation provided in (22) and (2.1), let us infer the statistical temporal structure of the velocity field $\mathbf{u}^{\nu}$ in the asymptotic limit $\nu\to 0$. The Hölder regularity pinpointed in (22) and (2.1) suggests that eddies of characteristic length scale $|\ell|$ have a characteristic turnover time scale $\tau_{\text{e}}(|\ell|)$ given by
$\tau_{\text{e}}(|\ell|)\equiv\frac{|\ell|}{\sqrt{S_{2}(\ell)}}\mathrel{\mathop{\kern 0.0pt\propto}\limits_{|\ell|\to 0}}\frac{|\ell|^{2/3}}{\sqrt{C_{2}}\varepsilon^{1/3}}.$ (28)
The square-root of the structure function $\sqrt{S_{2}(\ell)}$ can be interpreted as the typical velocity of eddies of size $|\ell|$. The characteristic turnover time $\tau_{\text{e}}(|\ell|)$ (28) of eddies of typical size $\ell$ would then follow a power-law with exponent $2/3$ as a function of the length scale $|\ell|$. One may wonder whether this prediction is correct. The proposed definition of the time scale $\tau_{\text{e}}(|\ell|)$ (28) clearly depends solely on the spatial structure of the velocity field and is independent of the temporal structure of the velocity field. In order to take into account the temporal structure, let us thus define a characteristic time scale $\tau_{c}(|\ell|)$ which is based on the correlation in time of the spatial increment given by
$\tau_{\text{c}}(|\ell|)\equiv\frac{1}{S_{2}(\ell)}\lim_{\nu\to 0}\int_{0}^{\infty}\mathbb{E}\left[\delta_{\ell}\mathbf{u}^{\nu}(t,\mathbf{x})\cdot\delta_{\ell}\mathbf{u}^{\nu}(t+\tau,\mathbf{x})\right]{\mathrm{d}}\tau.$ (29)
As it will be argued below, we could also define a similar time scale as $\tau_{\text{c}}(|\ell|)$ (29) to characterize the turnover time scale of an eddy of size $\ell$ using instead the temporal structure of the velocity Fourier modes at a given wave vector $k$. This indicates that one assimilates in a loose sense the length scale $|\ell|$ to $|\mathbf{k}|^{-1}$. This must be treated with great care since we have assumed in this Chapter that the spatial domain is infinite, i.e. $\mathbf{x}\in\mathbb{R}^{3}$, and thus Fourier modes must be considered as distributions. Nevertheless, we would define in a formal way the following time scale
@@
Note that the time scale $\tau_{\text{f}}(k)$ is expected to depend solely on the norm $k$ of the wave vector $\mathbf{k}$ by statistical isotropy and can be interpreted as the typical correlation duration at a given wave number. The two time scales $\tau_{\text{c}}(|\ell|)$ (29) and $\tau_{\text{f}}(k)$ (30) are related according to
$\tau_{\text{c}}(|\ell|)=``\frac{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left[1-\cos\left(2\pi\mathbf{k}\cdot\ell\right)\right]\tau_{\text{f}}(k)\mathbb{E}\left[\left|\widehat{\mathbf{u}}(t,\mathbf{k})\right|^{2}\right]{\mathrm{d}}\mathbf{k}}{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left[1-\cos\left(2\pi\mathbf{k}\cdot\ell\right)\right]\mathbb{E}\left[\left|\widehat{\mathbf{u}}(t,\mathbf{k})\right|^{2}\right]{\mathrm{d}}\mathbf{k}}$", (31)
@@
The behavior of the time scale $\tau_{\text{f}}(k)$ (32) at large wave numbers as $k^{-1}$ suggests, based on the dimensional ground, that the typical velocity entering in the motion of eddies of size $k^{-1}$ is the standard deviation of the velocity $\sigma$, i.e. the characteristic velocity of large scales. To see this more clearly, let us see the implications of this behavior on the time scale $\tau_{\text{c}}(|\ell|)$ (29). To do so, consider the identity displayed in (31) and rescale the dummy integration variable $\mathbf{k}$ by $|\ell|$, then one obtains
@@
$\tau_{\text{c}}(|\ell|)$ $=``\frac{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left[1-\cos\left(2\pi\mathbf{k}\cdot\frac{\ell}{|\ell|}\right)\right]\tau_{\text{f}}(|\mathbf{k}|/|\ell|)\mathbb{E}\left[\left|\widehat{\mathbf{u}}(t,\mathbf{k}/|\ell|)\right|^{2}\right]{\mathrm{d}}\mathbf{k}}{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left[1-\cos\left(2\pi\mathbf{k}\cdot\frac{\ell}{|\ell|}\right)\right]\mathbb{E}\left[\left|\widehat{\mathbf{u}}(t,\mathbf{k}/|\ell|)\right|^{2}\right]{\mathrm{d}}\mathbf{k}}$"
@@
$\mathrel{\mathop{\kern 0.0pt\propto}\limits_{|\ell|\to 0}}\frac{|\ell|}{\sigma}\frac{\int_{\rho\geq 0,\theta\in[0,\pi]}\left[1-\cos\left(2\pi\rho\cos\theta\right)\right]\rho^{-8/3}{\mathrm{d}}\rho{\mathrm{d}}\theta}{\int_{\rho\geq 0,\theta\in[0,\pi]}\left[1-\cos\left(2\pi\rho\cos\theta\right)\right]\rho^{-5/3}{\mathrm{d}}\rho{\mathrm{d}}\theta},$ (33)
@@
where we have used the asymptotic behaviors of $\tau_{\text{f}}(k)$ (32) and of the PSD $\mathbb{E}\left[\left|\widehat{\mathbf{u}}(t,\mathbf{k})\right|^{2}\right]\sim|\mathbf{k}|^{-11/3}$ (13) at large argument. Notice that the remaining ratio of double integrals entering in (2.3) can be shown to be finite (an exact expression can even be obtained using a symbolic calculation software). The relevant result given in (2.3) is the fact that, as a consequence of the observation that $\tau_{\text{f}}(k)$ becomes inversely proportional to $k$ at large $k$ (32), the genuine characteristic time scale $\tau_{\text{c}}(|\ell|)$ of eddies of typical size $|\ell|$ (29) becomes proportional to $|\ell|$ (2.3). Consequently, the relevant characteristic velocity appearing in (2.3) is $\sigma$, i.e., the characteristic velocity of the largest eddies, while it may have been more natural to obtain their typical velocity $\sqrt{S_{2}(\ell)}$ as in the expression (28) of $\tau_{\text{e}}(|\ell|)$: this is the phenomenon of advection of the small scales by the large eddies, referred to here and in the turbulence literature as the sweeping effect.
2.4 Implications for of the spatio-temporal covariance matrix of the velocity field
@@
Let us now present the general form of the covariance structure in time and space of the velocity field that encompasses both the spatial structure detailed in Paragraph 2.1 and the very peculiar temporal structure which is governed by the sweeping effect, as is in the former discussion 2.3. To do so, let us begin by defining the relevant spatio-temporal covariance of the velocity field $\mathcal{C}_{ij}^{\nu}(\tau,\ell)$ as
$\mathcal{C}_{ij}^{\nu}(\tau,\ell)=\mathbb{E}\left[u_{i}^{\nu}(t,\mathbf{x})u_{j}^{\nu}(t+\tau,\mathbf{x}+\ell)\right].$ (34)
@@
Again, $\mathcal{C}_{ij}^{\nu}(\tau,\ell)$ is independent of position $\mathbf{x}$ and time $t$ by statistical homogeneity and stationarity, respectively. We mention that, moreover, the spatiotemporal covariance function (34) is an even function of the temporal argument $\tau\in\mathbb{R}$, $\mathcal{C}_{ij}^{\nu}(-\tau,\ell)=\mathcal{C}_{ij}^{\nu}(\tau,\ell)$. In the following, we will also assume statistical isotropy and explore the consequences of this assumption on the structure of the covariance function. The natural extension of the spatial covariance matrix (6) to a spatio-temporal framework $\mathcal{C}_{ij}^{\nu}(\tau,\ell)$ (34) can be conveniently written in Fourier space as
@@
$\mathcal{C}_{ij}^{\nu}(\tau,\ell)=\frac{1}{2}\int_{\mathbf{k}\in\mathbb{R}^{3}}e^{2i\pi\mathbf{k}\cdot\ell}F(\tau/T_{\mathbf{k}})E_{\nu}^{\text{\tiny{E}}}(\mathbf{k})\widehat{P}_{ij}(\mathbf{k})\ {\mathrm{d}}\mathbf{k},$ (35)
@@
As we have already mentioned, it is clear from (35) that the PSD $E_{\nu}^{\text{\tiny{E}}}(k)$ corresponds to the Fourier transform of $\mathcal{C}_{ii}^{\nu}(0,\ell)$. As a consequence, we recover from (35) the scaling behaviors of the second-order structure functions (22) and (2.1). To make a connection with the discussion developed in Section 2.3, let us establish the link between the characteristic time scale $\tau_{\text{c}}(|\ell|)$ (29) and the $k$-dependent time scale $T_{k}$ entering in $\mathcal{C}_{ij}^{\nu}(\tau,\ell)$ (35). We have
$\tau_{\text{c}}(|\ell|)$ $=\frac{\int_{\tau=0}^{\infty}\left[\mathcal{C}_{ii}^{\nu}(\tau,0)-\mathcal{C}_{ii}^{\nu}(\tau,\ell)\right]{\mathrm{d}}\tau}{\mathcal{C}_{ii}^{\nu}(0,0)-\mathcal{C}_{ii}^{\nu}(0,\ell)}$
$=\frac{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\ell}\right)T_{\mathbf{k}}E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}}{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\ell}\right)E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}}\int_{0}^{\infty}F(s){\mathrm{d}}s.$ (36)
@@
where $D_{3}$ is a free parameter of the description that has the dimension of a velocity. Similarly to the derivation of the relationship between $\tau_{\text{c}}(|\ell|)$ and $\tau_{\text{f}}(k)$ obtained in (2.3), using the scaling behavior obtained in (37) and the behavior of the PSD in the same limit (13), we get from (2.4), while performing the remaining ratio of integrals expressed in spherical variables,
$\tau_{\text{c}}(|\ell|)$ $\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}\frac{|\ell|}{D_{3}}\frac{2^{1/3}\sqrt{\pi}\Gamma^{2}(1/6)}{15\Gamma(11/6)}\int_{0}^{\infty}F(s){\mathrm{d}}s.$ (38)
@@
Although physically insightful, the behavior of the characteristic time scale $\tau_{\text{c}}(|\ell|)$ with the eddy’s size $|\ell|$ (38) is not traditionally estimated in experiments and numerical simulations. Instead, the temporal second-order structure function and the time spectrum are usually studied.
@@
3 Gaussian random velocity vector, and its causal evolution

### v2

@@ 1 Introduction
$\mathbb{E}\left[\left(\delta_{\ell}^{\text{\tiny{long}}}\mathbf{u}^{\nu}\right)^{2}\right]\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}2\ell^{2}\int_{k\in\mathbb{R}}\left(2\pi k\right)^{2}E^{\text{\tiny{E,long}}}_{\nu}(k){\mathrm{d}}k\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\nu\to 0}}\frac{\varepsilon_{\nu}}{15\nu}\ell^{2},$ (27)
@@ 2.2 A first comparison concerning the spatial structure of DNS and the Gaussian model
We show in Fig. 1 a comparison between the velocity field $\mathbf{u}^{\text{\tiny{DNS}},\nu}(t,\mathbf{x})$ (Fig. 1(a)) extracted from a direct numerical simulations (DNSs) of the forced Navier-Stokes equations (1) and a Gaussian model (Fig. 1(b)). Concerning the DNS, it is provided by the Johns Hopkins turbulence database (Li et al., 2008). The data set concerns a simulation of the Navier-Stokes equations in a homogeneous and isotropic situation, using a pseudo-spectral method fully de-aliased with $N=1024$ nodes in each direction, over 5028 timesteps, corresponding to roughly speaking ten turnover time scales. The Taylor-based Reynolds number is estimated to be 433, corresponding to a fully developed turbulent state. Detailed properties of the simulation are provided on their website. For our purpose, we have downloaded the three velocity components over three spatial slices : $(0,y,z)$, $(x,0,z)$ and $(x,y,0)$, at all available times. The Gaussian model $\mathbf{u}^{\text{\tiny{fGv}},\nu}(\mathbf{x})$ corresponds to a fractional Gaussian vector field, in a periodic setting, presented in subsection 3.2.1. At this stage, it is enough to state that this is the unique statistically homogeneous, isotropic and incompressible Gaussian vector field consistent with the longitudinal PSD $E_{\nu}^{\text{\tiny{E,long}}}(k)$ provided in (10). Following a suitable truncation of the Fourier modes, as described in Section 4.1, we have similarly performed a simulation of this Gaussian vector field on a periodic box of size $L_{\text{\tiny{tot}}}=2\pi$ using $N=1024$ nodes in each direction, with the remaining free parameters entering in (10) chosen to match the DNS dataset. This corresponds to $D_{2}=0.021$, $L=L_{\text{\tiny{tot}}}$ and $\eta_{d}=0.085$. The results of our comparison are collected in Fig. 1.
@@
The statistical analysis of these two vector fields is shown in panels 1(c) and 1(d). The respective longitudinal PSDs $E_{\nu}^{\text{\tiny{E,long}}}(k)$ are shown in (c). Concerning the DNS, it is obtained as the variance of the respective Fourier modes obtained by performing one-dimensional discrete Fourier transforms of the longitudinal components, say, for instance, the $x$ component of velocity along the direction $x$. The statistical sample is averaged over the three longitudinal components, spatially over the slices, and over time. Results of our estimation are displayed with an orange solid line (light gray). The estimation of the Gaussian model statistics is also performed by averaging over the three components, over space and time, as has been done for the DNS. We display the result of our estimation using a solid blue line (dark gray). As we can observe, both longitudinal PSDs superimpose in a very satisfactory manner, showing that, based on the variance of Fourier modes, these two fields are statistically almost indiscernible. We mention that we could have, instead of the estimate on the Gaussian model, displayed the theoretical functional form provided in (10) without noticeable difference, as expected. The result of our statistical estimation of the longitudinal structure function $S_{2,\nu}^{\text{\tiny{long}}}(\ell)$ (20) is displayed in Fig. 1(d), using an orange line (light gray) for DNS and a blue line (dark gray) for the Gaussian model, averages have been taken on the same statistical sample as the one considered for Fig. 1(c). The three ranges of scales, including decorrelation length scales of order of the size of the box $L_{\text{\tiny{tot}}}$, inertial and dissipative length scales, are reproduced in a very satisfactory manner by the model, as expected from the behaviors of the PSDs of Fig. 1(c).
@@ 2.3 The statistical temporal structure of the turbulent velocity field and the sweeping effect
From the statistical spatial structure of the velocity field formerly presented, in large part given by the scaling relation provided in (22) and (2.1), let us infer the statistical temporal structure of the velocity field $\mathbf{u}^{\nu}$ in the asymptotic limit $\nu\to 0$. The Hölder regularity pinpointed in (22) and (2.1) suggests that eddies of characteristic length scale $\ell$ have a characteristic turnover time scale $\tau_{\text{e}}(\ell)$ given by
$\tau_{\text{e}}(\ell)\equiv\frac{\ell}{\sqrt{S_{2}(\ell)}}\mathrel{\mathop{\kern 0.0pt\propto}\limits_{\ell\to 0}}\frac{\ell^{2/3}}{\sqrt{C_{2}}\varepsilon^{1/3}}.$ (28)
The square-root of the structure function $\sqrt{S_{2}(\ell)}$ can be interpreted as the typical velocity of eddies of size $\ell$. The characteristic turnover time $\tau_{\text{e}}(\ell)$ (28) of eddies of typical size $\ell$ would then follow a power-law with exponent $2/3$ as a function of the length scale $\ell$. One may wonder whether this prediction is correct. The proposed definition of the time scale $\tau_{\text{e}}(\ell)$ (28) clearly depends solely on the spatial structure of the velocity field and is independent of the temporal structure of the velocity field. In order to take into account the temporal structure, let us thus define a characteristic time scale $\tau_{c}(\ell)$ which is based on the correlation in time of the spatial increment given by
$\tau_{\text{c}}(\ell)\equiv\frac{1}{S_{2}(\ell)}\lim_{\nu\to 0}\int_{0}^{\infty}\mathbb{E}\left[\delta_{\ell}\mathbf{u}^{\nu}(t,\mathbf{x})\cdot\delta_{\ell}\mathbf{u}^{\nu}(t+\tau,\mathbf{x})\right]{\mathrm{d}}\tau.$ (29)
As it will be argued below, we could also define a similar time scale as $\tau_{\text{c}}(\ell)$ (29) to characterize the turnover time scale of an eddy of size $\ell$ using instead the temporal structure of the velocity Fourier modes at a given wave vector $k$. This indicates that one associates in a loose sense the length scale $\ell$ to $|\mathbf{k}|^{-1}$. This must be treated with great care since we have assumed in this Chapter that the spatial domain is infinite, i.e. $\mathbf{x}\in\mathbb{R}^{3}$, and thus Fourier modes must be considered as distributions. Nevertheless, we would define in a formal way the following time scale
@@
Note that the time scale $\tau_{\text{f}}(k)$ is expected to depend solely on the norm $k$ of the wave vector $\mathbf{k}$ by statistical isotropy and can be interpreted as the typical correlation duration at a given wave number. The two time scales $\tau_{\text{c}}(\ell)$ (29) and $\tau_{\text{f}}(k)$ (30) are related according to
$\tau_{\text{c}}(\ell)=``\frac{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left[1-\cos\left(2\pi\mathbf{k}\cdot\boldsymbol{\ell}\right)\right]\tau_{\text{f}}(k)\mathbb{E}\left[\left|\widehat{\mathbf{u}}(t,\mathbf{k})\right|^{2}\right]{\mathrm{d}}\mathbf{k}}{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left[1-\cos\left(2\pi\mathbf{k}\cdot\boldsymbol{\ell}\right)\right]\mathbb{E}\left[\left|\widehat{\mathbf{u}}(t,\mathbf{k})\right|^{2}\right]{\mathrm{d}}\mathbf{k}}$", (31)
@@
The behavior of the time scale $\tau_{\text{f}}(k)$ (32) at large wave numbers as $k^{-1}$ suggests, based on the dimensional ground, that the typical velocity entering in the motion of eddies of size $k^{-1}$ is the standard deviation of the velocity $\sigma$, i.e. the characteristic velocity of large scales. To see this more clearly, let us see the implications of this behavior on the time scale $\tau_{\text{c}}(\ell)$ (29). To do so, consider the identity displayed in (31) and rescale the dummy integration variable $\mathbf{k}$ by $\ell$, then one obtains
@@
$\tau_{\text{c}}(\ell)$ $=``\frac{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left[1-\cos\left(2\pi\mathbf{k}\cdot\frac{\boldsymbol{\ell}}{\ell}\right)\right]\tau_{\text{f}}(|\mathbf{k}|/\ell)\mathbb{E}\left[\left|\widehat{\mathbf{u}}(t,\mathbf{k}/\ell)\right|^{2}\right]{\mathrm{d}}\mathbf{k}}{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left[1-\cos\left(2\pi\mathbf{k}\cdot\frac{\boldsymbol{\ell}}{\ell}\right)\right]\mathbb{E}\left[\left|\widehat{\mathbf{u}}(t,\mathbf{k}/\ell)\right|^{2}\right]{\mathrm{d}}\mathbf{k}}$"
@@
$\mathrel{\mathop{\kern 0.0pt\propto}\limits_{\ell\to 0}}\frac{\ell}{\sigma}\frac{\int_{\rho\geq 0,\theta\in[0,\pi]}\left[1-\cos\left(2\pi\rho\cos\theta\right)\right]\rho^{-8/3}{\mathrm{d}}\rho{\mathrm{d}}\theta}{\int_{\rho\geq 0,\theta\in[0,\pi]}\left[1-\cos\left(2\pi\rho\cos\theta\right)\right]\rho^{-5/3}{\mathrm{d}}\rho{\mathrm{d}}\theta},$ (33)
@@
where we have used the asymptotic behaviors of $\tau_{\text{f}}(k)$ (32) and of the PSD $\mathbb{E}\left[\left|\widehat{\mathbf{u}}(t,\mathbf{k})\right|^{2}\right]\sim|\mathbf{k}|^{-11/3}$ (13) at large argument. Notice that the remaining ratio of double integrals entering in (2.3) can be shown to be finite (an exact expression can even be obtained using a symbolic calculation software). The relevant result given in (2.3) is the fact that, as a consequence of the observation that $\tau_{\text{f}}(k)$ becomes inversely proportional to $k$ at large $k$ (32), the genuine characteristic time scale $\tau_{\text{c}}(\ell)$ of eddies of typical size $\ell$ (29) becomes proportional to $\ell$ (2.3). Consequently, the relevant characteristic velocity appearing in (2.3) is $\sigma$, i.e., the characteristic velocity of the largest eddies, while it may have been more natural to obtain their typical velocity $\sqrt{S_{2}(\ell)}$ as in the expression (28) of $\tau_{\text{e}}(\ell)$: this is the phenomenon of advection of the small scales by the large eddies, referred to here and in the turbulence literature as the sweeping effect.
2.4 Implications for the spatio-temporal covariance matrix of the velocity field
@@
Let us now present the general form of the covariance structure in time and space of the velocity field that encompasses both the spatial structure detailed in Paragraph 2.1 and the peculiar temporal structure which is governed by the sweeping effect, as is in the former discussion 2.3. To do so, let us begin by defining the relevant spatio-temporal covariance of the velocity field $\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})$ as
$\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})=\mathbb{E}\left[u_{i}^{\nu}(t,\mathbf{x})u_{j}^{\nu}(t+\tau,\mathbf{x}+\boldsymbol{\ell})\right].$ (34)
@@
Again, $\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})$ is independent of position $\mathbf{x}$ and time $t$ by statistical homogeneity and stationarity, respectively. We mention that, moreover, the spatiotemporal covariance function (34) is an even function of the temporal argument $\tau\in\mathbb{R}$, $\mathcal{C}_{ij}^{\nu}(-\tau,\boldsymbol{\ell})=\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})$. In the following, we will also assume statistical isotropy and explore the consequences of this assumption on the structure of the covariance function. The natural extension of the spatial covariance matrix (6) to a spatio-temporal framework $\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})$ (34) can be conveniently written in Fourier space as
@@
$\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})=\frac{1}{2}\int_{\mathbf{k}\in\mathbb{R}^{3}}e^{2i\pi\mathbf{k}\cdot\boldsymbol{\ell}}F(\tau/T_{\mathbf{k}})E_{\nu}^{\text{\tiny{E}}}(\mathbf{k})\widehat{P}_{ij}(\mathbf{k})\ {\mathrm{d}}\mathbf{k},$ (35)
@@
As we have already mentioned, it is clear from (35) that the PSD $E_{\nu}^{\text{\tiny{E}}}(k)$ corresponds to the Fourier transform of $\mathcal{C}_{ii}^{\nu}(0,\ell)$. As a consequence, we recover from (35) the scaling behaviors of the second-order structure functions (22) and (2.1). To make a connection with the discussion developed in Section 2.3, let us establish the link between the characteristic time scale $\tau_{\text{c}}(\ell)$ (29) and the $k$-dependent time scale $T_{k}$ entering in $\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})$ (35). We have
$\tau_{\text{c}}(\ell)$ $=\frac{\int_{\tau=0}^{\infty}\left[\mathcal{C}_{ii}^{\nu}(\tau,0)-\mathcal{C}_{ii}^{\nu}(\tau,\boldsymbol{\ell})\right]{\mathrm{d}}\tau}{\mathcal{C}_{ii}^{\nu}(0,0)-\mathcal{C}_{ii}^{\nu}(0,\boldsymbol{\ell})}$
$=\frac{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\boldsymbol{\ell}}\right)T_{\mathbf{k}}E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}}{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\boldsymbol{\ell}}\right)E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}}\int_{0}^{\infty}F(s){\mathrm{d}}s.$ (36)
@@
where $D_{3}$ is a free parameter of the description that has the dimension of a velocity. Similarly to the derivation of the relationship between $\tau_{\text{c}}(\ell)$ and $\tau_{\text{f}}(k)$ obtained in (2.3), using the scaling behavior obtained in (37) and the behavior of the PSD in the same limit (13), we get from (2.4), while performing the remaining ratio of integrals expressed in spherical variables,
$\tau_{\text{c}}(\ell)$ $\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}\frac{\ell}{D_{3}}\frac{2^{1/3}\sqrt{\pi}\Gamma^{2}(1/6)}{15\Gamma(11/6)}\int_{0}^{\infty}F(s){\mathrm{d}}s.$ (38)
@@
Although physically insightful, the behavior of the characteristic time scale $\tau_{\text{c}}(\ell)$ with the eddy’s size $\ell$ (38) is not traditionally estimated in experiments and numerical simulations. Instead, the temporal second-order structure function and the time spectrum are usually studied.
@@
3 Gaussian random velocity vector, and its Markovian evolution